User name gilman
Log entry time 16:02:38 on July31,2003
Entry number 108243
keyword=day shift summary
Day shift summary for 31 July 2003
Plan for shift: optics studies, Moller measurement, elastic scattering at 2.8 GeV
8:00 waiting for beam
9:35 MCC asks about sending CW beam into hall and tuning it up.
9:50 run harp scan on beam, horiz width ~ 160, vertical width ~ 490 (sigma)
10:00 beam off for beam studies
10:30 Folts et al into Hall on controlled access to work on OSHA
preparations, call him again if there are more chances on later day shifts.
12:15 Do controlled-access checklist, call for beam.
12:55 Beam back and roughyl centereed, start HARP scan, run 2310,
right arm DAQ run 3061. Apparent widths about 145 u horiz by 400 u vert.
13:08 ask MCC to reduce Y width by 2.
13:20 MCC has adjusted beam and asks for another HARP scan.
Scan run 2311 gives about 142 u by 410 u. Ooops.
13:24 ask for another HARP. Now 138 u vert by 419 u horiz.
13:32 Sirish directs operator to adjust "Compton quad". Run 2313 gives
46 u in X, 138 u in Y.
13:40 HARP scan 2314 gives 60 u by 180 u for scanner 6. While this is going
on, JP applies magic to fix left arm DAQ, which had hung. (Maybe not
magic, but he refuses to tell us what he did.)
13:45 HARP scan 2315, scanner 6 gives 57 u by 158 u.
13:53 HARP scan 2316, scanner 6 gives 53 u by 164 u.
14:03 HARP scan 2316, scanner 6 gives 53 u by 125 u.
Following JP's advice, we up threshold on target otr, and
now it gives more stable width reading. We then talk to MCC about the otr.
??:?? HARP scan 2317, scanner 6 gives 53 u by 125 u.
??:?? HARP scan 2318, scanner 6 gives 368 u by 311 u.
15:00 HARP scan 2319,
MCC now working on getting good setup through Compton chicane.
15:59 HARP scan 2320, scanner 6 gives 382 u by 310 u
With Compton people grabbing power, we are changing over to the next shift!